can you help me by solving completing the square?
x^2-2x-35=0

Answer by dolly(163)

x^2 - 2x - 35 = 0
adding 35 to both the sides
==> x^2 - 2x = 35
add 1 to both the sides,
x^2 - 2x + 1 = 35 + 1
==> (x-1)^2 = 36
Taking square root,
x - 1 = +/- 6
==> x-1 = 6 or x-1 = -6
==> x = 7 0r x = -5
Thus x = 7 or x = -5